Lynch Company manufactures and sells a single product. The following costs were incurred during the company’s first year of operations:

Variable costs per unit:
Manufacturing:
Direct materials $ 10
Direct labor $ 4
Variable manufacturing overhead $ 2
Variable selling and administrative $ 2
Fixed costs per year:
Fixed manufacturing overhead $ 374,000
Fixed selling and administrative $ 284,000

During the year, the company produced 34,000 units and sold 26,000 units. The selling price of the company’s product is $44 per unit.

Required:

1. Assume that the company uses absorption costing:

a. Compute the unit product cost.

b. Prepare an income statement for the year.

2. Assume that the company uses variable costing:

a. Compute the unit product cost.

b. Prepare an income statement for the year.
Answer #1

Answer-1-a)- Unit product cost under Absorption costing:-Direct materials + Direct Labor+ Variable manufacturing overhead + fixed manufacturing overhead

=$10+$4+$2+$11 = $27 per unit

Unit fixed manufacturing overhead= fixed manufacturing overhead/No. of units produced

=$374000/34000 units =$11 per unit

b)-

Lynch Company
Income statement (Using absorption costing approach)
Particulars Amount
$
Sales (a) 26000 units*$44 per unit 1144000
Less:- Cost of goods sold (b)
Opening inventory
Add:- Cost of goods manufactured 918000
Direct materials 34000 units*$10 per unit 340000
Direct labor 34000 units*$4 per unit 136000
Variable manufacturing overhead 34000 units*$2 per unit 68000
Fixed manufacturing overhead 374000
Cost of goods available for sale 918000
Less:- Closing inventory 8000 units*$27 per unit 216000 702000
Gross margin C= a-b 442000
Less:-Variable selling & administrative exp. 26000 units*$2 per unit 52000
Less:- Fixed costs
Selling & administrative exp. 284000
Net Income 106000

Answer-2-a)-Unit product cost under Variable costing:-Direct materials + Direct Labor+ Variable manufacturing overhead

=$10+$4+$2 = $16 per unit

2-b)-

Lynch Company
Income statement (Using variable costing approach)
Particulars Amount
$
Sales (a) 26000 units*$44 per unit 1144000
Less:- Variable cost of goods sold (b)
Opening inventory NIL
Add:- Variable cost of goods manufactured 544000
Direct materials 34000 units*$10 per unit 340000
Direct labor 34000 units*$4 per unit 136000
Variable manufacturing overhead 34000 units*$2 per unit 68000
Variable cost of goods available for sale 544000
Less:- Closing inventory 8000 units*$16 per unit 128000 416000
Gross contribution margin C= a-b 728000
Less:-Variable selling & administrative exp. 26000 units*$2 per unit 52000
Contribution margin 676000
Less:- Fixed costs
Manufacturing overhead 374000
Selling & administrative exp. 284000
Net Income 18000